OpenStudy (anonymous):

-5x - 3y = -16 -x + 9y = 64 elimination method

OpenStudy (ajprincess):

Multiply the second eqn by -5 and tell me vat u get. Can u?

OpenStudy (anonymous):

5x-45y=320

OpenStudy (anonymous):

Good! Now add it to the first one. Please tell us what do you get.

OpenStudy (ajprincess):

Vat do u get?

OpenStudy (anonymous):

x-48y=48

OpenStudy (ajprincess):

Can u try again? -5x - 3y = -16-(1) 5x-45y=320-(2) Add (1) and (2) Nw plz tell vat u get?

OpenStudy (jiteshmeghwal9):

First try to make one of the variable equal in both of the equation like\[2(-5x-3y)=(-16)2\]\[10(-x+9y)=(64)10\]